Rosa invests $17,000 at 4% simple interest for 41 years.
Round your answers to the nearest cent.
How much interest is earned over the 41 year period?
$
How much is in the account at the end of the 41 year period?
$

Answer by jhunjiro(67)

A = P (1 + rt)
A = final amount
P = initial principal balance
r = annual interest rate
t = time (in years)
Given
A = ?
P = 17000
r = 4%
t = 41 years
Solution:
A = 17000(1 + (4%)(41))
A = 17000(1 + 1.64)
A = 17000(2.64)
A = 44,880
How much interest is earned over the 41 year period?
interest is earned over the 41 year period = 44,880 - 17000
-
interest is earned over the 41 year period = $27,880
-
How much is in the account at the end of the 41 year period?
the account at the end of the 41 year period = $44,880